OpenStudy (anonymous):

If 12b = 36, what is the value of b?

OpenStudy (alexandervonhumboldt2):

just divide both sides by 12: 12b/12=36/12 b=3

OpenStudy (anonymous):

b = 3 b = 4 b = 18 b = 24